# A Code Knowledge Graph-Enhanced System for LLM-Based Fuzz Driver Generation

### v1

@@ IV-C1 Code knowledge graph
To evaluate the impact of the code knowledge graph on the quality of fuzz drivers generated by CodeGraphGPT, we compared it with a variant that uses text-based API knowledge from libraries. This variant generates API combinations by querying a text-based knowledge base consisting of API source code and API summaries. The experimental results in Table II show that, the text-only retrieval variant achieves lower code coverage compared to CodeGraphGPT, this highlights the effectiveness of the code knowledge graph in generating high-quality fuzz drivers.
@@ IV-C2 Compilation repair
In CodeGraphGPT, we introduced a dynamic program repair component to fix fuzz driver programs that encounter compilation errors. In this section, we design experiments to evaluate the effectiveness of CodeGraphGPT in repairing compilation errors and the impact of the dynamic program repair component on fuzzing performance. First, we created two variants: one without any compilation error repair (without repair) and another that relies solely on an LLM for program repair (LLM-only repair). The experimental results in Table II show that the without repair achieved an average compilation success rate of only 57.39% (458/798), indicating that CodeGraphGPT struggles to directly generate compilable fuzz driver programs. On the other hand, the LLM-only repair increased the compilation success rate to 77.19% (616/798), demonstrating the significant potential of LLMs for program repair. However, CodeGraphGPT achieved a compilation success rate of 93.99% (750/798), highlighting the effectiveness of the dynamic program repair component in fixing compilation errors for fuzz driver programs.
Additionally, we evaluated the code coverage performance of the two variants. The results in Table II show that both the without repair and LLM-only repair achieved lower code coverage compared to CodeGraphGPT. This is due to many fuzz drivers containing unique API combinations that failed to compile and, thus, could not participate in the fuzzing process, limiting the potential to explore new execution paths in the libraries through diverse API combinations.
Figure 2: Branch coverage of CodeGraphGPT with and without coverage-guided mutation in different libraries
@@ IV-C3 Coverage-guided mutation
CodeGraphGPT utilizes coverage-guided mutation to discover potential new API combinations. To assess its effectiveness, we conducted a comparative experiment against a variant that does not apply API combination mutation. We kept the fuzzing time fixed at 24 hours for both approaches to compare their coverage trends. Figure 2 shows the branch coverage results of CodeGraphGPT with and without API combination mutation across different libraries.
Our coverage-guided mutation demonstrates a performance advantage across all eight libraries, although the benefit is less pronounced in cjson and libvpx. This is because the number of APIs in these two libraries is relatively small and functionally similar, meaning that nearly all possible API combinations were explored during the initial API combination generation phase. In the remaining six libraries, coverage-guided mutation significantly improved the code coverage of CodeGraphGPT. The experimental results suggest that coverage-guided mutation is a more effective approach in most cases.
@@
TABLE III: Summary of crash analysis results in CodeGraphGPT
@@
CodeGraphGPT incorporates a crash analysis module to automate the detection of crash causes during fuzzing. As shown in Table III, we collected the results from this crash analysis module.
@@ IV-D1 Overall result
CodeGraphGPT identified 199 unique crashes during fuzzing, of which 168 were determined by the crash analysis module to be caused by the misuse of library APIs. To minimize false positives, the crash analysis module treats cases that are ambiguous or difficult to classify as misuse crashes. For example, if a crash is caused by passing a non-existent or error pointer to a library API, and the API does not employ defensive programming to check the validity of the input pointer, we still categorize this as a misuse crash. To verify the accuracy of the crash analysis module in detecting misuse crashes, we randomly sampled 10% of the misuse crash cases from each library for manual inspection. After reviewing these cases, we confirmed that all of them were indeed caused by the misuse of the API by the fuzz driver.
@@ IV-D2 Case study
Figure 3: Bugs found by CodeGraphGPT
3(a) shows a buffer overflow bug case discovered by CodeGraphGPT. In this code, the unsafe strcpy function is used without checking the length of the input string f. Since the data passed by the fuzz driver is randomly generated, the length of f is not controlled, and it may exceed the size of the it8->DoubleFormatter buffer. This leads to a buffer overflow, which is a severe security vulnerability that could potentially be exploited by attackers to overwrite memory and execute arbitrary code.
3(b) illustrates another bug case discovered by CodeGraphGPT, which is an out-of-memory (OOM) issue. In this case, *tif refers to a TIFF file controlled by CodeGraphGPT. The library function does not properly validate the fields such as width and length within the file. As a result, it returns an abnormal tile value. Other library APIs attempt to allocate memory based on the calculated tile size, which causes excessive memory allocation, ultimately leading to an OOM crash. Our case study demonstrates the effectiveness of CodeGraphGPT in detecting real-world bugs. We will publicly disclose the bug cases we have discovered on our website.
@@ V-A External
The evaluation of CodeGraphGPT was conducted using a set of eight widely-used open-source libraries. While these libraries cover a range of API usage patterns and functionalities, the generalizability of our approach to other libraries or domains remains uncertain. It is possible that CodeGraphGPT’s performance may vary when applied to libraries with different structures, API designs, or use cases. Another potential threat is bias introduced by the training data of the LLM used in CodeGraphGPT. The LLM may favor over-represented patterns in its training set, which could affect the generation of API combinations or the crash analysis process. To mitigate this risk, we employed diverse prompts and knowledge sources to guide the LLM and reduce the impact of any biases.
@@ V-B Internal
A key internal threat to validity is the potential for LLM hallucinations during code generation. Incorrect or irrelevant code snippets may be generated, requiring multiple rounds of refinement or manual intervention to correct. This could affect the quality and functionality of the generated fuzz drivers. Additionally, the performance of CodeGraphGPT depends heavily on the accuracy and completeness of the knowledge graph and API information extracted from the codebase. Incomplete or inaccurate knowledge may result in suboptimal API combinations or ineffective crash analysis.
@@ VI-A Applicability to Different Programming Languages
CodeGraphGPT was primarily evaluated on C-based libraries, which are widely used and have well-documented APIs. However, the applicability of CodeGraphGPT to other programming languages, such as Rust, Java, or Python, remains an open question. Since each language has different API designs, CodeGraphGPT may require adjustments to effectively handle these differences. Future work could explore the adaptation of CodeGraphGPT to these languages, potentially broadening its scope and utility.
@@ VI-B Quality of the Code Knowledge Graph
The code knowledge graph plays a central role in CodeGraphGPT by guiding API combination generation and updating API combinations during coverage-guided mutation. The completeness and accuracy of the code knowledge graph are essential for generating effective fuzz drivers. In our experiments, libraries with more comprehensive code knowledge graph showed better coverage and higher fuzzing effectiveness. Improving the automation, construction, and validation of the code knowledge graph could further enhance the performance of CodeGraphGPT by ensuring that it accurately reflects the library’s API structure and usage patterns.

### v2

@@ IV-C1 Code knowledge graph
To evaluate the impact of the code knowledge graph on the quality of fuzz drivers generated by CKGFuzzer, we compared it with a variant that uses text-based API knowledge from libraries. This variant generates API combinations by querying a text-based knowledge base consisting of API source code and API summaries. The experimental results in Table II show that, the text-only retrieval variant achieves lower code coverage compared to CKGFuzzer, this highlights the effectiveness of the code knowledge graph in generating high-quality fuzz drivers.
@@ IV-C2 Compilation repair
In CKGFuzzer, we introduced a dynamic program repair component to fix fuzz driver programs that encounter compilation errors. In this section, we design experiments to evaluate the effectiveness of CKGFuzzer in repairing compilation errors and the impact of the dynamic program repair component on fuzzing performance. First, we created two variants: one without any compilation error repair (without repair) and another that relies solely on an LLM for program repair (LLM-only repair). The experimental results in Table II show that the without repair achieved an average compilation success rate of only 57.39% (458/798), indicating that CKGFuzzer struggles to directly generate compilable fuzz driver programs. On the other hand, the LLM-only repair increased the compilation success rate to 77.19% (616/798), demonstrating the significant potential of LLMs for program repair. However, CKGFuzzer achieved a compilation success rate of 93.99% (750/798), highlighting the effectiveness of the dynamic program repair component in fixing compilation errors for fuzz driver programs.
Figure 2: Branch coverage of CKGFuzzer with and without coverage-guided mutation in different libraries
Additionally, we evaluated the code coverage performance of the two variants. The results in Table II show that both the without repair and LLM-only repair achieved lower code coverage compared to CKGFuzzer. This is due to many fuzz drivers containing unique API combinations that failed to compile and, thus, could not participate in the fuzzing process, limiting the potential to explore new execution paths in the libraries through diverse API combinations.
@@ IV-C3 Coverage-guided mutation
CKGFuzzer utilizes coverage-guided mutation to discover potential new API combinations. To assess its effectiveness, we conducted a comparative experiment against a variant that does not apply API combination mutation. We kept the fuzzing time fixed at 24 hours for both approaches to compare their coverage trends. Figure 2 shows the branch coverage results of CKGFuzzer with and without API combination mutation across different libraries.
Our coverage-guided mutation demonstrates a performance advantage across all eight libraries, although the benefit is less pronounced in cjson and libvpx. This is because the number of APIs in these two libraries is relatively small and functionally similar, meaning that nearly all possible API combinations were explored during the initial API combination generation phase. In the remaining six libraries, coverage-guided mutation significantly improved the code coverage of CKGFuzzer. The experimental results suggest that coverage-guided mutation is a more effective approach in most cases.
@@
TABLE III: Summary of crash analysis results in CKGFuzzer
@@
CKGFuzzer incorporates a crash analysis module to automate the detection of crash causes during fuzzing. As shown in Table III, we collected the results from this crash analysis module.
@@ IV-D1 Overall result
CKGFuzzer identified 199 unique crashes during fuzzing, of which 168 were determined by the crash analysis module to be caused by the misuse of library APIs. To minimize false positives, the crash analysis module treats cases that are ambiguous or difficult to classify as misuse crashes. For example, if a crash is caused by passing a non-existent or error pointer to a library API, and the API does not employ defensive programming to check the validity of the input pointer, we still categorize this as a misuse crash. To verify the accuracy of the crash analysis module in detecting misuse crashes, we randomly sampled 10% of the misuse crash cases from each library for manual inspection. After reviewing these cases, we confirmed that all of them were indeed caused by the misuse of the API by the fuzz driver.
@@ IV-D2 Case study
Figure 3: Bugs found by CKGFuzzer
3(a) shows a buffer overflow bug case discovered by CKGFuzzer. In this code, the unsafe strcpy function is used without checking the length of the input string f. Since the data passed by the fuzz driver is randomly generated, the length of f is not controlled, and it may exceed the size of the it8->DoubleFormatter buffer. This leads to a buffer overflow, which is a severe security vulnerability that could potentially be exploited by attackers to overwrite memory and execute arbitrary code.
3(b) illustrates another bug case discovered by CKGFuzzer, which is an out-of-memory (OOM) issue. In this case, *tif refers to a TIFF file controlled by CKGFuzzer. The library function does not properly validate the fields such as width and length within the file. As a result, it returns an abnormal tile value. Other library APIs attempt to allocate memory based on the calculated tile size, which causes excessive memory allocation, ultimately leading to an OOM crash. Our case study demonstrates the effectiveness of CKGFuzzer in detecting real-world bugs. We will publicly disclose the bug cases we have discovered on our website.
@@ V-A External
The evaluation of CKGFuzzer was conducted using a set of eight widely-used open-source libraries. While these libraries cover a range of API usage patterns and functionalities, the generalizability of our approach to other libraries or domains remains uncertain. It is possible that CKGFuzzer’s performance may vary when applied to libraries with different structures, API designs, or use cases. Another potential threat is bias introduced by the training data of the LLM used in CKGFuzzer. The LLM may favor over-represented patterns in its training set, which could affect the generation of API combinations or the crash analysis process. To mitigate this risk, we employed diverse prompts and knowledge sources to guide the LLM and reduce the impact of any biases.
@@ V-B Internal
A key internal threat to validity is the potential for LLM hallucinations during code generation. Incorrect or irrelevant code snippets may be generated, requiring multiple rounds of refinement or manual intervention to correct. This could affect the quality and functionality of the generated fuzz drivers. Additionally, the performance of CKGFuzzer depends heavily on the accuracy and completeness of the knowledge graph and API information extracted from the codebase. Incomplete or inaccurate knowledge may result in suboptimal API combinations or ineffective crash analysis.
@@ VI-A Applicability to Different Programming Languages
CKGFuzzer was primarily evaluated on C-based libraries, which are widely used and have well-documented APIs. However, the applicability of CKGFuzzer to other programming languages, such as Rust, Java, or Python, remains an open question. Since each language has different API designs, CKGFuzzer may require adjustments to effectively handle these differences. Future work could explore the adaptation of CKGFuzzer to these languages, potentially broadening its scope and utility.
@@ VI-B Quality of the Code Knowledge Graph
The code knowledge graph plays a central role in CKGFuzzer by guiding API combination generation and updating API combinations during coverage-guided mutation. The completeness and accuracy of the code knowledge graph are essential for generating effective fuzz drivers. In our experiments, libraries with more comprehensive code knowledge graph showed better coverage and higher fuzzing effectiveness. Improving the automation, construction, and validation of the code knowledge graph could further enhance the performance of CKGFuzzer by ensuring that it accurately reflects the library’s API structure and usage patterns.